Part 1 Timely Transport provides local delivery service for a number of downtown and suburban businesses. Delivery charges are based on distance and weight involved for each delivery: 20 cents per pou...

Part 1

Timely Transport provides local delivery service for a number of downtown and suburban businesses. Delivery charges are based on distance and weight involved for each delivery: 20 cents per pound and 15 cents per mile. Also, there is a $25 handling fee per parcel.

a. Develop an expression that summarizes delivery charges. (Enter your answers as dollar amounts to 2 decimal places (x1 = weight, x2 = distance). Omit the "$" sign in your response.)

y = $ _____ x1 + $_____ x2 +$ ______

b. Determine the delivery charge for transporting a 35-pound parcel 15 miles. (Do not round your intermediate calculation. Round your answer to 2 decimal places. Omit the "$" sign in your response.)

y = $ ______

Part 2

The manager of a fashionable restaurant open Wednesday through Saturday says that the restaurant does about 38 percent of its business on Friday night, 22 percent on Saturday night, and 12 percent on Thursday night. What seasonal relatives would describe this situation?(Round your answers to 2 decimal places.)

Wednesday
Thursday
Friday
Saturday
Answer #1

Part 1 :

a. y = $25 + $ (0.2) x1 + (0.15) x2

b. Substituing the values,

25 + 0.2*35 + 0.15*15 = $34.25

Part 2:

Wednesday = 0.28*4 = 1.12

Thursday = 0.12*4 = 0.48

Friday = 0.38*4 = 1.52

Saturday = 0.22*4 = 0.88

PLEASE LEAVE A LIKE IF YOU FIND THIS HELPFUL. THANKS A LOT.